

********************************************************************
UARS WEEKLY STATUS REPORT - 30 JULY 1999
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to operate normally, collecting solar data on all
available orbital tracking periods and stellar calibration observations,
whenever possible.
***********
SUSIM
***********
SUSIM resumed operations on 7/26. Subsequently, it gathered its normal
solar, offset, and occultation data without apparent problems.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
The PEM spectrometers continue to be active under PMON control. While
under PMON control, the PEM instrument computed energy deposition from its
particle spectrometers.
The FOT and the PEM team are completing close-out documentation on
the HEPS Energy Processor trouble; however progress has been hampered
with the departure of Peter Thomas.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 17 July 1999 (Orbit 42878, GMT Day 198) through 23 July
(GMT Day 204).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing) and
is performing nominally under two-battery operations. The instrument
operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
WINDII 200/1206 42915 On
PEM 202/2140 42951 AXIS Off
ACRIM 202/2140 42951 Off
SUSIM 202/2140 42951 Off
WINDII 202/2140 42951 Off
MLS 203/0443 42955 On
HRDI 204/2102 42977 Reset
The beta angle ranged between the angles of 21.8 degrees to 50.4 degrees
for this report period. The beta angle is now increasing toward a maximum
angle of 75 degrees on 31 July 1999 (DOY 212).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage.
Battery temperatures are stable with a temperature delta between Battery 2
and Battery 3 of 0.9 to 1.1 degrees C. Additional battery performance data
for each battery on 16 July 1999 and 23 July 1999 are provided in Appendix
D.
The clock error ranged from 11.4 to -11.2 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT for
this report. A summary is listed below and the details of ALERT and Out-
Of-Limits occurrences are listed in Appendix B.
% ALERTS = 00
% OUT-OF-LIMITS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
% FOT 03
% WINDII 02
% HRDI 01
% PEM 01
% ACRIM 01
% SUSIM 01
% MLS 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
% FOT 00
One (1) Anomaly Report remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
Transponder A: 99/179 - 2287.496941
Transponder B: 99/180 - 2287.498977
GROUND SYSTEM OPERATION
Zero (0) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from
mission start through UTC Day 99/203 show the observatory data loss to be
24 hours, 54 minutes, 38.644 seconds (an increase of 20.576 seconds since
the last report period). This is a 0.0363 percent data loss which equals a
99.9637 percent data capture for the mission.
The increase of 20.576 seconds occurred over three (3) day:
7 8.736 seconds on Day 198 due to DRC not covered,
7 0.640 seconds on Day 201 due to DRC not covered,
7 11.200 seconds on Day 202 due to no overlap.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
FOT personnel are supporting the IMOC (SCS-21) development effort. The
current work is a continuation of the SCS-21 UARS implementation effort
done last year which culminated in a level 1 demonstration.
The FOT continues to investigate the feasibility of a new PMON action to
change battery V/T curves based on loading to mitigate battery risks in sun
pointing safehold.
YEAR 2000 OPERABILITY SUPPORT
The FOT received a set of Y2K planning aids from FDF which will be used to
facilitate UARS Y2K end-to-end testing. Eleven 3-day planning aid files
were received and processed by the FOT. Two of these files did not process
successful on the CMS due to either incorrect data, or invalid data
formats. FDF has been notified of the problems, and are in the process of
correcting the files.
MEETINGS
FOT attended a meeting with GRO. FOT to discuss their recent anomalies and
lessons learned for applications to UARS.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw Around Maneuver (forward-to reverse) is scheduled for 20
August 1999 (DOY 232).
PERSONNEL CHANGES
None
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 23 July 1999
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
42912 200/0655 99-243 OBC Clk Adj 0.00 to 20.6 msec/day
42915 200/1206 99-241 WINDII Power ON
42938 202/0059 99-244 OBC Clk Adj 20.6 to 0.00 msec/day
42951 202/2140 99-245 AXIS Power OFF
42951 202/2140 99-246 ACRIM Power OFF
42951 202/2140 99-247 SUSIM Power OFF
42951 202/2140 99-248 WINDII Power OFF
42955 203/0443 99-249 MLS Power ON
42970 204/0448 99-250 OBC Clk Adj 0.00 to 20.6 msec/day
42977 204/2102 99-251 HRDI HRDI RESET
_______________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 23 July 1999
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
None this report period
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
None this report period
_______________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 23 July 1999
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
42912 200/07:03 (19 July 99) from 00.0 to 20.6 msec/day
42938 202/01:09 (20 July 99) from 20.6 to 00.0 msec/day
42970 204/04:53 (23 July 99) from 00.0 to 20.6 msec/day
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 23 July 1999
16 July 1999 - GMT Day 197 Beta = 21.8 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 -100.8/-123.2 -3.05/-4.31 +0.0/ +0.0 16.5 0.000 0.0
2 +44.8/-28.0 +5.79/+4.83 +30.4/-14.4 25.0 1.023 21.8
3 +28.0/-33.6 +4.83/+3.87 +29.6/-14.8 24.8 1.023 21.5
23 July 1999 - GMT Day 204 Beta = 50.4 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------ ------
1 -106.4/-134.4 -3.05/-4.62 +0.0/ +0.0 16.5 0.000 0.0
2 +44.8/-22.4 +5.47/+4.19 +26.4/-16.0 25.1 1.021 21.0
3 +22.4/-50.4 +4.51/+3.23 +25.6/-15.6 25.0 1.023 20.7
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- -------- ----------
no changes V/T 5
APPENDIX E
UARS ANOMALY REPORTS
Week ending 23 July 1999
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
None this report period
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
None this report period
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------- ------ ----- ---- --- -----------------
None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-033 FOT see list July 99 14170 Generic Late Acq.
SEVNTFW99/30